# ST-Booster: An Iterative SpatioTemporal Perception Booster for Vision-and-Language Navigation in Continuous Environments

### v1

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) requires agents to navigate unknown, continuous spaces based on natural language instructions. Compared to discrete settings, VLN-CE poses two core perception challenges. First, the absence of predefined observation points leads to heterogeneous visual memories and weakened global spatial correlations. Second, cumulative reconstruction errors in three-dimensional scenes introduce structural noise, impairing local feature perception. To address these challenges, this paper proposes ST-Booster, an iterative spatiotemporal booster that enhances navigation performance through multi-granularity perception and instruction-aware reasoning. ST-Booster consists of three key modules — Hierarchical SpatioTemporal Encoding (HSTE), Multi-Granularity Aligned Fusion (MGAF), and Value-Guided Waypoint Generation (VGWG). HSTE encodes long-term global memory using topological graphs and captures short-term local details via grid maps. MGAF aligns these dual-map representations with instructions through geometry-aware knowledge fusion. The resulting representations are iteratively refined through pretraining tasks. During reasoning, VGWG generates Guided Attention Heatmaps (GAHs) to explicitly model environment-instruction relevance and optimize waypoint selection. Extensive comparative experiments and performance analyses are conducted, demonstrating that ST-Booster outperforms existing state-of-the-art methods, particularly in complex, disturbance-prone environments.
Index Terms:
Vision-and-language navigation, robot navigation, map learning.
{NTP}
Vision-and-Language Navigation in Continuous Environments demands robust perception to navigate reconstructed real-world scenes under observational noise. This paper presents ST-Booster, a practical solution that enhances scene understanding by combining global layout tracking with local detail mapping. ST-Booster maintains fused spatiotemporal environmental representations by integrating navigation memory with real-time fine-grained observations, which are further aligned with linguisitc instructions. It autonomously identifies instruction-relevant regions via predicted heatmaps to inform path planning decisions. This multi-scale perception strategy significantly improves navigation robustness in noisy, cluttered indoor environments. To support real-world applications, we open-source the code of ST-Booster, enabling its deployment in indoor autonomous navigation tasks for robotics practitioners.
@@ I Introduction
Vision-and-Language Navigation (VLN) is an interdisciplinary indoor embodied task [1, 2, 3] that seeks to develop intelligent robots to navigate complex environments based on natural language instructions. Traditionally, VLN frameworks operate within discrete environments, where robots move only between predefined waypoints. In contrast, VLN in continuous environments (VLN-CE) [4, 5], trains robots to navigate smoothly in continuous spaces through small, incremental actions. Although VLN-CE offers a more realistic scenario for navigation tasks, it also introduces substantial challenges. Firstly, the absence of the predefined navigation graphs compels agents to establish long-term spatiotemporal dependencies through diverse visual observations at arbitrary positions. Secondly, the inherent visual noise in continuous environments during three-dimensional (3D) reconstruction severely compromises local observation accuracy. These factors collectively necessitate enhanced spatiotemporal perception capabilities for robust visual interpretation and trajectory-instruction alignment.
Figure 1: An overview of environmental representation methods in VLN-CE and the proposed ST-Booster framework. (a) Topological map-based representations for global historical perceptions [6]. (b) Grid map-based representations for local spatial environments [7]. (c) The proposed ST-Booster framework for iterative enhancement with hybrid maps. (d) Pretraining tasks utilized in ST-Booster for robust action reasoning.
@@
Current approaches for VLN-CE are broadly categorized into end-to-end methods and the map-based methods. End-to-end methods implicitly learn environmental perceptions through RGB-D image processing [8, 9], facing challenges in preserving fine-grained spatial details and long-term history due to feature compression. Map-based methods have emerged as solutions for achieving effective environmental perception, owing to their inherent advantages in structured representation [10, 11, 12]. Particularly, as shown in Figure 1 (a), topological mapping techniques [6, 13, 14, 15] address long-term dependencies by constructing graph representations, however, they often oversimplify the local representations. Conversely, grid-based methods [7, 16] in Figure 1 (b) excel in capturing detailed local features but face the challenge of computational efficiency for large-scale environments. Recent hybrid strategies attempt to synergize these complementary map representations. Initial efforts [17] directly combined two different map features through averaging operations, which diluted critical spatial relationships. Subsequent advancements [18] introduced multi-scale fusion mechanisms. However, the prevailing late-fusion paradigm inherits limitations from individual map representations.
@@
Figure 2: The proposed ST-Booster comprises three core modules. First, the Hierarchical SpatioTemporal Encoding (HSTE) Module captures global spatial structures via global topological graphs and extracts local temporal details using local grid-based maps. Next, the Multi-Granularity Aligned Fusion (MGAF) Module fuses these heterogeneous map features and aligns the fused representations with linguistic embeddings. Finally, the integrated representation are utilized in the Value Guided Waypoint Generation (VGWG) module to predict Guided Attention Heatmaps (GAHs), which adaptively adjust candidate waypoint distributions to instruction-relevant regions.
To address these perception challenges in VLN-CE, we propose ST-Booster, an iterative spatiotemporal enhancement method with three core components illustrated in Figure 2. First, the Hierarchical SpatioTemporal Encoding (HSTE) module decouples long- and short-term perception through complementary representations. Global topological graphs aggregate historical trajectory patterns to establish structured global memory, while local grid maps employ deformable projections to capture fine-grained environmental details. Second, Multi-Granularity Aligned Fusion (MGAF) module bridges these heterogeneous representations via bidirectional cross-domain fusion. Specifically, topological nodes are projected onto grid coordinates with spatial distribution weights to inject historical context into local features, while local details from grid maps are embedded into corresponding topological nodes through region matching to enhance global precision. Cross-modal alignment with linguistic instructions further integrates textual intent into the fused representations. Finally, this multi-granularity perception undergoes iterative optimization through several pretraining tasks, including a weakly-supervised Guided Attention Heatmap (GAH) prediction task that explicitly models environment-instruction joint probabilities through a Value-Guided Waypoint Generation (VGWG) module. VGWG predicts GAHs and utilizes them to guide the generated candidate waypoints to prioritize instruction-relevant regions, overcoming the limitation of conventional waypoint generation methods that ignore linguistic constraints. Experiments demonstrate improved navigation success rates in complex scenarios with environmental interference, compared to state-of-the-art methods evaluated on VLN-CE datasets.
The remainder of this paper is organized as follows. Section II conducts a comprehensive review of the relevant literature. The design of ST-Booster is presented in Section III. Section IV details dataset construction and presents the experimental results. Finally, Section V provides the concluding remarks.
@@ II Related Work
Vision-and-Language Navigation in Continuous Environments (VLN-CE). Facing the challenges of VLN-CE, researchers have explored various techniques. Initial methods for VLN-CE attempted to predict low-level actions from continuous visual observations through end-to-end training [4, 19]. However, these methods failed to effectively process historical information due to the extended navigation paths with overlapping observations. To address this limitation, researchers introduced waypoint predictors, which leverage predicted local navigation graphs to provide candidate waypoints for subsequent decision-making [20, 21, 22]. This modular approach, inspired by discrete VLN, bridges the gap between discrete and continuous environments and has been further extended with semantic information [23, 24, 25]. Nevertheless, the performance of VLN-CE is still far below VLN [26, 27]. This mainly arises from the fact that both end-to-end and modular approaches rely on implicit visual representations derived from complex inputs, hindering the understanding of spatiotemporal relationships in environments and diminishing the model’s capacity to capture the vision-textual dependencies. To overcome these challenges, this paper proposes a novel environmental perception booster that combines the structural advantages of global topological and local grid maps to explicitly model spatial-temporal relationships.
@@
Environment Representation. Scene perception is essential for indoor navigation, offering agents visual cues and contextual insights [28, 29, 30, 31]. Traditional end-to-end training methods often lack structured constraints, leading to loss and distortion of representation. To address this issue, researchers have explored map-based approaches to effectively manage perceptual data during navigation [32, 33, 34]. For instance, Chen et al. [6] introduced a framework combining topological maps and attention mechanisms for navigation planning. In a later study [35], an online topological map is utilized to enhance the agent’s spatial comprehension. Some other approaches leverage grid maps to represent environmental details. Unlike topological graphs that compress visual representations into nodes, grid maps offer detailed local spatial information, aiding short-term decision-making [36, 37]. For instance, Georgios et al. [23] proposed a cross-modal map learning framework to enhance path prediction. Meanwhile, Huang et al. [32] integrated pre-trained vision-language features with 3D reconstruction for spatial mapping. In order to combine the strengths of topological and grid maps, hybrid map methods have been proposed. For instance, Liu et al. [17] constructed a global topological map by averaging grid map features and fused the decision results. In contrast, An et al. [18] enhanced both maps by incorporating panoramic image features and their low-dimensional representations, and improved action reasoning by combining independent decisions. However, existing hybrid map methods are still constrained by the inherent limitations of each map type without feature-level fusion. This paper proposes an iterative spatiotemporal perception booster that focuses on feature-level hybrid map fusion to achieve a more comprehensive perception and improve the model’s multi-granularity understanding of the environment.
@@ III ST-Booster
In this section, we first outline the problem formulation of VLN-CE (Section III-A). Subsequently, we introduce ST-Booster, an iteratively spatiotemporal booster to enhance agent perception in complex environments. As illustrated in Figure 2, HSTE module constructs hybrid maps by introducing global topological and local grid map representations (Section III-B). These hybrid map representations are fused and aligned with instructions through MGAF module (Section III-C). The obtained multi-granularity features are utilized in VGWG module to predict GAHs, which optimize the candidate waypoint generation with instruction-relevant knowledge (Section III-D). The multi-granularity features are iteratively refined using several pretraining tasks and then fine-tuned with online-generated maps for navigation reasoning (Section III-E).
@@ III-A Problem Formulation
[…]
@@ III-B Hierarchical SpatioTemporal Encoding Module
To effectively navigate complex continuous environments, we propose a hierarchical spatiotemporal encoding module that decomposes environmental representation into long-term and short-term components.
@@
Structural Topology Encoding for Global Perception. At step $t$, the global topological map $\mathcal{G}_{t}=\{\mathcal{V}_{t},\mathcal{E}_{t}\}$ captures long-term environmental awareness by jointly modeling visited and observed waypoints. The node set $\mathcal{V}_{t}=\mathcal{V}^{\text{v}}_{t}\cup\mathcal{V}^{\text{o}}_{t}$ stores the visited nodes $\mathbf{n}\in\mathcal{V}^{\text{v}}_{t}$ with averaged panoramic visual representations and the observed nodes $\mathring{\mathbf{n}}\in\mathcal{V}^{\text{o}}_{t}$ with view-specific visual representations, where $\mathcal{V}^{\text{v}}_{t}=\{\mathbf{n}_{i}\}_{i=1}^{t}$ and $\mathcal{V}^{\text{o}}_{t}=\{\{\mathring{\mathbf{n}}_{i}^{k}\}_{k=1}^{g_{i}}\}_{i=1}^{t}$, $g_{i}$ is the number of observed waypoints at $i-$th time step. In particular, the visual representations include the RGB features extracted via ViT [39] and the depth features extracted from ResNet [40]. In addition, the edge set $\mathcal{E}_{t}$ encodes the Euclidean distances between adjacent nodes. The nodes in $\mathbf{G}_{t}$ dynamically integrate perception context through fused relative position embeddings and time-step embeddings. Additionally, a globally connected virtual node $\mathbf{n}_{0}$ is created to represent the ‘stop’ action. Topological maps efficiently manage global scaling and maintain historical memory, but compress the representation details. To address this, we integrate local grid maps to enrich fine-grained environmental representation.
Figure 3: Illustration of our feature fusion approach: visual observations are embedded in topological nodes and grid cells. Node features are flattened by global spatial relations to add historical context to grid features, while grid aggregation supplies fine-grained spatial details to nodes.
Dense Grid Encoding for Local Perception. To represent the detailed surrounding environment, we introduce a local grid map $\mathbf{M}_{t}\in\mathbb{R}^{U\times U\times D}$ centered on the agent’s current position, where $U$ and $D$ denote the map scale size and feature dimensions, respectively. To mitigate computational complexity, we construct $\mathbf{M}_{t}$ using dimensionality-reduced RGB features and depth images. Leveraging the agent’s pose and depth data, the downsized RGB features are projected onto the ground grid through the spatial transformation between 2D and 3D coordinate spaces. Similar to node embeddings, for each grid cell, $\mathbf{m}_{u,v}$, in addition to introducing the projected features, we also incorporate position embeddings to capture the spacial relationships between the grid cells. Grid maps offer precise local information but are limited in scalability and historical context.

### v2

@@ Abstract
Vision-and-Language Navigation in Continuous Environments (VLN-CE) requires agents to navigate previously unseen and continuous spaces based on natural language instructions. Compared to discrete settings, VLN-CE poses two core perception challenges. First, the absence of predefined observation points leads to heterogeneous visual memories and weakened global spatial correlations. Second, cumulative reconstruction errors in three-dimensional scenes introduce structural noise, impairing local feature perception. To address these challenges, this paper proposes ST-Booster, an iterative spatiotemporal booster that enhances navigation performance through multi-granularity perception and instruction-aware reasoning. ST-Booster consists of three key modules — Hierarchical SpatioTemporal Encoding (HSTE), Multi-Granularity Aligned Fusion (MGAF), and Value-Guided Waypoint Generation (VGWG). HSTE encodes long-term global memory using topological graphs and captures short-term local details via grid maps. MGAF aligns these dual-map representations with instructions through geometry-aware knowledge fusion. The resulting representations are iteratively refined through pretraining tasks. During reasoning, VGWG generates Guided Attention Heatmaps (GAHs) to explicitly model environment-instruction relevance and optimize waypoint selection. Extensive comparative experiments and performance analyses are conducted, demonstrating that ST-Booster outperforms existing state-of-the-art methods, particularly in complex, disturbance-prone environments. Our code is available at https://github.com/yueluhhxx/ST-Booster.
@@ I Introduction
Vision-and-Language Navigation (VLN) [anderson2018vision, tan2025source, wu2021improved] is an embodied intelligence task where agents follow natural-language instructions to navigate using visual inputs from limited data [wang2024vision, wang2023res]. Traditionally, VLN frameworks operate within discrete environments, where robots move only between predefined waypoints. In contrast, VLN in continuous environments (VLN-CE) [krantz2020beyond, anderson2021sim] requires low-level control in continuous space. Although VLN-CE offers a more realistic scenario for navigation tasks, it also introduces substantial challenges. As illustrated in Fig. 1(a), without a predefined graph structure, the agent reasons over long horizons and handles noisy observations from challenging viewpoints at arbitrary poses in three-dimensional (3D) reconstructed environments. This requires stronger spatiotemporal perception and instruction alignment.
Figure 1: Overview of environmental representation and perception in VLN-CE. (a) Comparison of observations between discrete environments (MP3D) and continuous environments (Habitat); (b) advantages and limitations of topological [chen2021topological] and grid maps [georgakis2022cross]; (c) workflow of BEVBert [an2023bevbert] using dual-map decision fusion; (d) workflow of ST-Booster with iterative dual-map feature fusion and multi-task decision integration.
@@
Current approaches for VLN-CE are broadly categorized into end-to-end and map-based methods. End-to-end methods implicitly learn environmental perceptions through RGB-D image processing [hong2021vln, chen2021history, hong2024navigating], facing challenges in preserving fine-grained spatial details and long-term history due to feature compression. Map-based methods as illustrated in Fig. 1(b) have emerged as solutions for achieving effective environmental perception, owing to their inherent advantages in structured representation. Particularly, topological mapping techniques [chen2021topological, huo2023geovln] address long-term dependencies by constructing graph representations; however, they often oversimplify the local representations. Conversely, grid-based methods [wang2023gridmm] excel in capturing detailed local features but suffer from high computational cost for large-scale environments. Recent hybrid strategies utilize both complementary map representations. Initial efforts [liu2023bird] directly combined two map features through averaging operations, which diluted critical spatial relationships. Subsequent advancements [an2023bevbert] introduced multi-scale fusion mechanisms (Fig. 1(c)). However, the prevailing late-fusion paradigm inherits limitations from individual map representations.
@@
Figure 2: The proposed ST-Booster comprises three core modules. First, the Hierarchical SpatioTemporal Encoding (HSTE) module captures global spatial structures via global topological graphs and extracts local temporal details using local grid-based maps. Next, the Multi-Granularity Aligned Fusion (MGAF) module fuses these heterogeneous map features and aligns the fused representations with linguistic embeddings. Finally, the integrated representations are utilized in the Value Guided Waypoint Generation (VGWG) to predict Guided Attention Heatmaps (GAHs), which adaptively adjust candidate waypoint distributions to instruction-relevant regions.
To tackle the perception challenges in VLN-CE, we introduce ST-Booster, an iterative spatiotemporal enhancement framework built upon the dual-map knowledge sharing idea (Fig. 1(d)). ST-Booster consists of three main modules (Fig. 2). First, the Hierarchical SpatioTemporal Encoding (HSTE) module balances long- and short-term spatiotemporal dependencies by constructing global topological maps for long-term reasoning and local grid maps for detailed perception. Although these maps are complementary, their heterogeneous representations hinder effective feature-level transfer. To address this, we propose the Multi-Granularity Aligned Fusion (MGAF) module, which fuses maps bidirectionally by adding global context to grids and local details to topological nodes for unified semantic understanding. Finally, to enhance interpretability and strengthen decision-level reasoning, we include the Value-Guided Waypoint Generation (VGWG) module, which predicts Guided Attention Heatmaps (GAHs) to highlight instruction-relevant regions and generate meaningful candidate waypoints. Extensive experiments demonstrate that ST-Booster achieves robust and consistent navigation in noisy and complex environments, effectively integrating spatial and temporal cues for more reliable perception and decision making.
The contributions of this paper are threefold. First, this paper introduces ST-Booster, an iterative refinement framework that integrates the advantages of global topological and local grid maps in perceiving spatiotemporal information at the feature space level. Second, this paper designs a weakly supervised module that predicts Guided Attention Heatmaps (GAHs) to focus candidate waypoints on instruction-relevant regions, improving decision robustness and interpretability. Third, extensive experiments on VLN-CE validate the effectiveness of the proposed framework, demonstrating consistent performance gains over state-of-the-art methods, particularly in unseen and noisy environments.
@@ II Related Work
Vision-and-Language Navigation in Continuous Environments (VLN-CE). Facing the challenges of VLN-CE, researchers have explored various techniques. Initial methods attempted to predict low-level actions from continuous visual observations through end-to-end training [krantz2020beyond, raychaudhuri2021language, zhao2023mind]. However, these methods struggled to capture long-term dependencies along paths with overlapping observations. To address this limitation, researchers introduced waypoint predictors, which leverage predicted local navigation graphs to provide candidate waypoints for subsequent decision-making [krantz2021waypoint, hong2022bridging, krantz2022sim]. This modular approach bridges the gap between discrete and continuous environments and has been extended with semantic information [wu2024vision, an2021neighbor, georgakis2022cross, chen2022weakly, irshad2021sasra]. Nevertheless, VLN-CE performance remains below discrete VLN [li2023improving, wang2023dreamwalker]. A key reason is that both end-to-end and modular approaches rely on implicit visual representations, hindering spatiotemporal reasoning and vision-textual dependencies. To enhance structured spatiotemporal reasoning, we extend the dual-map expert framework of An et al. [an2023bevbert] with an iterative spatiotemporal perception booster that performs feature-level fusion and cross-map knowledge transfer. This mechanism integrates complementary global and local cues into a complete scene representation. Furthermore, a multi-task joint training scheme reinforces the fused features and provides instruction-aware priors for sampling multimodally aligned waypoint candidates, improving both navigation accuracy and interpretability.
@@
Environment Representation. Scene perception is essential for navigation, offering agents visual cues and contextual insights [min2021film, duh2020v]. Traditional end-to-end training methods often lack structured constraints, leading to loss and distortion of representation. To address this issue, researchers have explored map-based approaches to effectively manage perceptual data [huang2023visual, jiang2024bevnav]. For instance, Chen et al. [chen2021topological] introduced a framework combining topological maps and attention for planning. In a later study [chen2024mapgpt], an online topological map further improves spatial understanding. Some other approaches leverage grid maps to represent environmental details. Unlike topological graphs that compress visual representations into nodes, grid maps offer detailed local spatial information, aiding short-term decision-making. For instance, Georgios et al. [georgakis2022cross] proposed a cross-modal map learning framework to enhance path prediction. Meanwhile, Huang et al. [huang2023visual] integrated pre-trained vision-language features with 3D reconstruction for spatial mapping. To combine their strengths, hybrid map methods have been proposed. For instance, Liu et al. [liu2023bird] constructed a global topological map by averaging grid map features and fused the decision results. In contrast, An et al. [an2023bevbert] enhanced both maps by incorporating panoramic image features and their low-dimensional representations, improving action reasoning via decision fusion. However, existing hybrid map methods are still constrained by the inherent limitations of each map type without feature-level fusion. This paper proposes an iterative spatiotemporal perception booster that focuses on feature-level hybrid map fusion to achieve more comprehensive, multi-granularity perception.
@@ III ST-Booster
In this section, we present the problem formulation of VLN-CE (Section III-A) and introduce ST-Booster, an iterative spatiotemporal framework for perception in complex environments. As illustrated in Fig. 2, the HSTE module builds hybrid maps by combining global topological and local grid representations (Section III-B), which are fused and aligned with instructions through the MGAF module (Section III-C). The resulting multi-granularity features are used by the VGWG module to predict GAHs for instruction-relevant waypoint generation (Section III-D) and are iteratively refined through pretraining and fine-tuning with online-generated maps for navigation reasoning (Section III-E).
@@ III-A Problem Formulation
Following the conventional setting of VLN-CE [an2024etpnav, an2023bevbert], an agent receives an instruction containing $L$ words for each navigation episode. This instruction is embedded as $\mathbf{I}=\{{\bm{l}_{1},\cdots,\bm{l}_{L}}\}$, where $\bm{l}_{i}$ represents the embedding of the $i$-th word. During the navigation process, the agent records its position information and observes the surrounding panoramic RGB-D images $\mathbf{O}_{t}=\{(\mathbf{o}^{\text{rgb}}_{t,i},\mathbf{o}^{\text{d}}_{t,i})\}_{i=0}^{11}$ at each time step $t$. These images are taken from 12 distinct views, each oriented at a horizontal heading angle of $\theta_{i}=30^{\circ}\times i$, where $i=0,\cdots,11$. Here, $\mathbf{o}^{\text{rgb}}_{t,i}\in\mathbb{R}^{H\times W\times 3}$ and $\mathbf{o}^{\text{d}}_{t,i}\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ represent the RGB and depth images observed from the $i$-th view, respectively. The agent’s objective is to effectively integrate textual and visual information to navigate toward the target locations specified in the instruction. In contrast to selecting discrete waypoints, agents in VLN-CE navigate within a 3D continuous environment by choosing actions from a low-level action space defined as $\mathscr{A}=\{\text{forward 0.25m},\text{turn left }15^{\circ},\text{turn right }15^{\circ},\text{stop}\}$.
@@ III-B Hierarchical SpatioTemporal Encoding Module
To effectively navigate complex continuous environments, we propose a hierarchical spatiotemporal encoding module that decomposes the environmental representation into long-term and short-term components.
@@
Structural Topology Encoding for Global Perception. At step $t$, the global topological map $\mathcal{G}_{t}=\{\mathcal{V}_{t},\mathcal{E}_{t}\}$ captures long-term environmental awareness by jointly modeling visited and observed waypoints. The node set $\mathcal{V}_{t}=\mathcal{V}^{\text{v}}_{t}\cup\mathcal{V}^{\text{o}}_{t}$ stores the visited nodes $\mathbf{n}\in\mathcal{V}^{\text{v}}_{t}$ with averaged panoramic visual representations and the observed nodes $\mathring{\mathbf{n}}\in\mathcal{V}^{\text{o}}_{t}$ with view-specific visual representations, where $\mathcal{V}^{\text{v}}_{t}=\{\mathbf{n}_{i}\}_{i=1}^{t}$ and $\mathcal{V}^{\text{o}}_{t}=\{\{\mathring{\mathbf{n}}_{i,k}\}_{k=1}^{g_{i}}\}_{i=1}^{t}$, $g_{i}$ is the number of observed waypoints at $i$-th time step. In particular, the visual representations include the RGB features extracted via ViT [dosovitskiy2020image] and the depth features extracted from ResNet [he2016deep]. In addition, the edge set $\mathcal{E}_{t}$ encodes the Euclidean distances between adjacent nodes. The nodes in $\mathbf{G}_{t}$ dynamically integrate perception context through fused relative position embeddings and time-step embeddings. Additionally, a globally connected virtual node $\mathbf{n}_{0}$ is created to represent the ‘stop’ action. Topological maps efficiently manage global scaling and maintain historical memory, but compress the representation details. To address this, we integrate local grid maps to enrich fine-grained environmental representation.
Dense Grid Encoding for Local Perception. To represent the detailed surroundings of the agent, we construct a local grid map $\mathbf{M}_{t}\in\mathbb{R}^{U\times V\times D}$ centered on the agent, where $U$ and $V$ are spatial dimensions and $D$ is the feature dimension. To reduce computational cost while maintaining geometric fidelity, $\mathbf{M}_{t}$ is built from dimensionality-reduced RGB features and aligned depth inputs. Specifically, patch-level RGB embeddings $\mathbf{V}^{\text{M-R}}_{t}\in\mathbb{R}^{12\times P\times P\times D}$ are extracted using a ViT encoder [dosovitskiy2020image] with patch size $P$. The corresponding depths ${\mathbf{o}^{\text{d}}_{t,i}},i=0,1,\ldots,11$ are average-pooled to the same $P\times P$ layout, denoted as $\mathbf{d}_{t}=\text{AvgPool}({\mathbf{o}^{\text{d}}_{t,i}})$. Each depth value in $\mathbf{d}_{t}$ is then back-projected into 3D and transformed into the ego-centric frame using the current pose, forming a point cloud $\mathbf{P}_{t}=[\mathbf{p}_{t,1},\dots,\mathbf{p}_{t,N}]$ with $N=12\times P\times P$ points. Each 3D point $\mathbf{p}_{t,i}$ inherits its RGB embedding $\mathbf{V}^{\text{M-R}}_{t,i}$, producing semantically enriched point features. To obtain a compact spatial representation, the point features are projected onto a Bird’s-Eye-View (BEV) grid of size $U\times V$. Invalid or out-of-range points are removed, and valid points are assigned to grid cells based on their coordinates. For each grid cell $\mathbf{m}_{(u,v)}$, the aggregated feature is computed as $\mathbf{m}_{(u,v)}=\frac{1}{|\mathcal{S}_{(u,v)}|}\sum_{i\in\mathcal{S}_{(u,v)}}\mathbf{V}^{\text{M-R}}_{t,i}$, where $\mathcal{S}_{(u,v)}$ denotes the set of points falling into that cell. This process converts the sparse 3D point cloud into a structured BEV feature map $\mathbf{M}_{t}$, providing a compact spatial representation for downstream reasoning. For each grid cell $\mathbf{m}_{(u,v)}$, positional embeddings are added to encode spatial relationships between neighboring cells. The resulting grid map offers fine-grained local environmental information, complementing global representations.